If u is the initial velocity of the body projected with an angle theta with the horizontalย  then maximum height reached is?Please answer in an easy way. Thanks ๐Ÿ™๐Ÿ™

Grade:9

2 Answers

Arun
4 years ago
Dear student
ย 
Maximum height reached = u^2 sin^2 theta / 2g
ย 
Hope it helps
ย 
Regards
Arun (askIITians forum expert)
Khimraj
4 years ago

H max.=(u^2 sin^2theta/2g) where u is initial velocity and g =9.8 m/s^2 and theta is the angle of projection.

If x=90 then it means it is thrown vertically and h max.=u^2/2g..